Problem 7-05 Nonconstant Growth Valuation A company currently pays a dividend of $3.25 per share (D0 = $3.25). It is estimated that the company's dividend will grow at a rate of 15% per year for the next 2 years, and then at a constant rate of 7% thereafter. The company's stock has a beta of 0.9, the risk-free rate is 5.5%, and the market risk premium is 5%. What is your estimate of the stock's current price? Do not round intermediate calculations. Round your answer to the nearest cent.
Answer #1

Required rate=risk free rate+Beta*market risk premium

=5.5+(0.9*5)=10%

D1=(3.25*1.15)=3.7375

D2=(3.7375*1.15)=4.298125

Value after year 2=(D2*Growth rate)/(Required rate-Growth rate)

=(4.298125*1.07)/(0.1-0.07)

=153.299792

Hence current price=Future dividend and value*Present value of discounting factor(rate%,time period)

=3.7375/1.1+4.298125/1.1^2+153.299792/1.1^2

=$133.64(Approx).